Contract a table to answer and draw a bar paragraph;


Ashipments of five computer contains two that are slightly defective. If a retailer receives three of these computers at random, list the elements of the same space S using the letters D and N for defective and non-defective computers, respectively. To each sample point, assign a value x of the random variable X representing the number of computers purchased by the retailer which are slightly defective


1
Expert's answer
2021-03-04T14:36:43-0500



"\text{Probability space:}\\
S=\{DDD, DDN,NDD,DND,DNN,NDN,NND,NNN\}\\
P\{X=0\}=(3/5)^3\\
P\{X=1\}=3(2/5)(3/5)^2\\
P\{X=2\}=3(2/5)^2(3/5)\\
P\{X=3\}=(2/5)^3"